Posts

NPP Projection Close 2/6

To do this evening/ was completed: A slower night on the projection side of things, despite the many guests in the building this evening. Lots of manual starts and calls for bringing up and down lights (which should continue into tomorrow with the number of overlaps between rounds)   Presentation Issues: Preshow Clips/Playlist: Trailer Updates: Special Events/ AV set up: Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Outgoing/Shipping (content): Print/video Inspection: Spent the majority of the evening finishing inspecting  Love Battlefield  for Sundays on Fire. Reels 2-5 did not have any leaders or cues so I added those. A couple of the leaders I used were taken off spare trailers from booth 1 and can be reattached after the screening Asset tracking doc updated/checked?: Maintenance/cleaning: Cue sheets updated/checked (every shift): Yes Marquee updated?: Posters updated?: Cleaning lights on for the night?: Yes Projection cart by entrance?: Ye...

NPP Projection Admin 2/6

To do this evening/ was completed: A1 2/13 playlists built, old content removed from servers. Tech'd Peter Hujar, all good. Tech'd Dead Lover too, all good, but we'll need to adjust masking presets to properly cover 1.33 in T3. Sent out a zoom link to the participants of the Q&A plus Lillian and Cristina. Cable tested in T2 for Super Bowl. Rescheduled T1 and T4 for today and tomorrow. Cue sheets updated. Dog Day Afternoon print arrived, in T1. Presentation Issues: Preshow Clips/Playlist: Trailer Updates: Special Events/ AV set up: Techd and adjusted the playlist for Blazing Saddles, which will be for the private event on Sunday in T6. - LH Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Ingested the 2D version of Goat , we had the 3D one on the LMS -A Ingested new content for the week (what was available) for preshow Sundays on Fire V4, Dog Day afternoon preshow, forecasts, etc. Waiting on a few other bits n bobs. - LH  Ingested Kubo and the T...

NPP Projection Close 2/5

To do this evening/ was completed: Presentation Issues: Theater 3's server froze during the Q&A and lost connection to the internet, forcing me to reboot the projector as the moderator and guests were finishing up their discussion, and delaying the next and final round by 5 minutes Preshow Clips/Playlist: Trailer Updates: Special Events/ AV set up: Removal of the Eye screening and Q&A in theater 3 went smoothly save for the aforementioned technical issue that kept me from unpausing the show playlist and starting the next round Alec got STB up and running as I was handling the Q&A in theater 3...I left the gear beneath the table cloth for the show on Saturday but disconnected the spotlight board to plug the masking motor back into the outlet (I made sure the table was out of the path of the curtain/tested scope and flat settings and it is functioning normally). Mics and laptop are back in the booths charging for the time being. Ordering/Receiving (posters & trailers)...

Projection Admin 2/5 Lillian (Alec on mid)

To do this evening/ was completed: There is a headset in T5 at seat 306. Please pick this up once the feature is over. Thanks! Presentation Issues: Preshow Clips/Playlist: Took a look at The Moment today, which I brought up to 5.5 as some dialogue can be soft, but the music is loud enough.  We're moving the Superbowl to T2, swapping the 6:45 The Moment and 9:15 Send Help shows to 3. Made this adjustment on the schedule and cue sheet for Sunday. -A Trailer Updates: Special Events/ AV set up: Setup for Spoons today in T1. Tested the wireless mics, laptop and spots. All you need to do is turn everything on, plus the breaker for the spots and bring down the mics from their charging ports once it is time. The mic stands are by the table, also set up.In addition to this, the three mics and chairs are all set for Removal of the Eye in T3, which I got eyes on just to double check. --Made sure they were all set before heading out. Tested mics, picture, sound, house and spot lights. All g...

NPP Projection Close 2/4

To do this evening/ was completed: T3 Q&A went smoothly, mics broken down afterward. NoBudge playlist built. One of the shipping reels Let The Right One In came with is broken so I didn't package that reel up yet. We have a spare in T1 if we want to use that? Put Dancing Lady and Jurassic Park on house reels. Zip file of Make Out With Violence downloaded to 4/6 desktop. Presentation Issues: Preshow Clips/Playlist: Trailer Updates: Special Events/ AV set up: Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Outgoing/Shipping (content): Print/video Inspection: Asset tracking doc updated/checked?: Maintenance/cleaning: Cue sheets updated/checked (every shift): Yes Marquee updated?: Posters updated?: Cleaning lights on for the night?: Projection cart by entrance?: Additional Notes:

Projection Admin 2/4

To do this evening/ was completed: Weekly update is out - please look things over as soon as you can and familiarise yourself with any pertinent info for shows you'll be working. Thanks! NOTE - forgot to mention this in the update, but the private event for Sirat will be at a later date (still holding, though) this month. We will be showing this as a first run anyhow, so it'll be on the server still. - LH  Presentation Issues: Preshow Clips/Playlist: Went through the Frankenhooker playlist/feature just to double check everything. All good. No intro/memo.  Trailer Updates: Special Events/ AV set up: Setup the mics and chairs (off to the side) in T3 for Monk in Pieces tonight. Just remember to turn on the mixer and you can put the mics on the chairs if you'd like. Mics have been tested. Friendly reminder - spots remote/power switch in 3/7 booth. I noticed there wasn't a slide for the Q&A in the playlist, so that's been adjusted. LH The direct line from the mic m...

NPP Projection Close 2/3

To do this evening/ was completed: Presentation Issues: Preshow Clips/Playlist: Trailer Updates: Special Events/ AV set up: Set up for trivia this afternoon and put back all equipment used following the show in theater 1.  Let the Right One In screening went well. All reels are tails out beneath the rewind bench in booth 1. Both projectors were cleaned after the show and the exhausts are shut for the night. Ordering/Receiving (posters & trailers): Dead Lover  posters were delivered and stored behind 3/7 Ordering/Receiving (content): Outgoing/Shipping (content): Print/video Inspection: Asset tracking doc updated/checked?: Maintenance/cleaning:  Cue sheets updated/checked (every shift): Yes Marquee updated?: Posters updated?: Cleaning lights on for the night?: Yes Projection cart by entrance?: Yes Additional Notes: Received a 3.5mm attenuation aux cable from Amazon and brought it up to 4/6